### v3

@@ Abstract
Unleashing the predictive power of molecular dynamics (MD), Neural Network Potentials (NNPs) trained on Density Functional Theory (DFT) calculations are revolutionizing our ability to simulate chemical systems with unprecedented accuracy and efficiency. Electrolyte solutions are a natural initial system to apply this tool to because they are critically important for a wide range of applications and their properties cannot currently be predicted. Unfortunately, however most DFT approximations are not sufficiently accurate to predict many practically relevant properties of electrolytes. Additionally, tracking the position of every atom in a system during molecular simulations is inherently limited, even with NNP-MD. Here, we use a state-of-the-art DFT approximation to demonstrate highly accurate all-atom NNPs with minimal training data. We demonstrate that NNPs can reliably be recursively trained on a subset of their own output to enable coarse-grained continuum solvent molecular simulations that can access much longer timescales. We apply our technique to simulate lithium chloride, potassium chloride, and lithium bromide in water. We reproduce key structural, thermodynamic, and kinetic properties of these solutions in agreement with experimental data. The formation of a previously unknown Li cation dimer is observed, along with identical anion-anion interactions of chloride and bromide. Finally, the coarse-grained model is capable of reproducing crystal phase behavior and infinite dilution pairing free energies despite being trained solely on moderate concentration solutions, disproving the notion that NNPs are only useful for interpolation. This approach should be scalable to determine the properties of electrolyte solutions over a much wider range of conditions and compositions than is possible experimentally.
@@ 1 Introduction
Molecular-scale processes, occurring at the level of thousands of atoms, are at the heart of chemistry and biology. These processes obey the laws of quantum mechanics and dictate the behavior of a vast range of crucial systems, yet they remain elusive, hidden from direct observation. We rely on indirect experiments and models to piece together their mysteries, but this limits our mastery over vital biological, chemical, and material systems. Unlocking these secrets could revolutionize our understanding and control of the world at its most fundamental level.
@@
Accurate and efficient first principles molecular dynamics simulations (FPMD) of these processes would be a transformatively useful tool for achieving this goal. This would enable the direct observation of key processes; the calculation of important properties using statistical mechanics; and the generation of abundant training data for machine learning models.
@@
This is now becoming possible thanks to neural network potentials (NNPs).[1, 2, 3, 4] This approach trains a neural network to reproduce quantum chemistry calculations, usually generated with density functional theory (DFT).[5, 6, 7, 8, 9, 10, 11, 12] NNPs are several orders of magnitude faster than direct calculations and are rapidly improving thanks to advances such as equivariance[13, 14, 15, 16, 17] and explicit electrostatics[18, 19, 7, 20, 21].
NNPs can also be used to run coarse-grained molecular dynamics simulations that ignore irrelevant parts of a systems, i.e., solvent, enabling additional orders of magnitude acceleration. These coarse-grained NNPs are normally trained on all-atom classical molecular dynamics.[22, 23, 24, 25, 26] This approach has yet to be demonstrated for more complex first principles potential energy surfaces.
Electrolyte solutions are a natural initial system to apply these tools to because determining the properties of electrolytes from first principles is a foundational problem of physical chemistry. These solutions play a key role in a vast range of important processes and systems. For example, lithium cations are the primary charge carriers for Li-ion batteries, and their chemical equilibria and diffusivities impact the assembly and performance of these devices.[27, 28] Lithium also exhibits important biochemical effects as a treatment for bipolar disorder. However, electrolytes are important in such a wide range of systems that highlighting individual cases does not properly convey the full scope of applications.
@@
Initial work has shown significant promise.[27, 29, 18, 30, 2, 31, 32, 33, 34, 31, 20, 35, 36, 37, 38, 21, 39, 40, 41, 42, 43, 44, 45, 46] However, quantitative prediction of important electrolyte solution properties such as activity coefficients that depend on ion pairing with reasonable training data requirements has yet to be demonstrated. This is a critical goal as chemical engineering of systems involving electrolytes currently relies on models fitted to empirical data, which significantly limits their applicability.[47] The main issue is that electrolyte solutions are both highly dynamic and have long-range electrostatic interactions. This has meant training NNPs for electrolytes has previously required large training datasets at several concentrations and active learning approaches. As a result, a relatively low level of DFT approximation has been used to generate the training data. This is a problem as many DFT approximations are not accurate enough to quantitatively predict electrolyte solution properties. [48, 49, 50, 51] It has also been particularly difficult to simulate ion-ion pairing at low concentrations because these pairs are much rarer than ion-solvent or solvent-solvent interactions.
@@
Figure 1: Workflow: Small, short first principles molecular dynamics (FPMD) simulations are run with CP2K. Energies/forces computed with DC-r2SCAN are used to train an neural network potential (NNP), which enables much faster simulations with a larger cell. Forces and coordinates of the ions alone are output from the NNP-MD and used to train a coarse-grained continuum solvent NNP, which enables even faster MD simulations and the simulation of crystal dissolution, despite being trained only on the solution-phase.
@@
Density-corrected DFT (DC-DFT) has been shown to reduce errors, such as delocalization and self-interaction, which cause standard density functional approximations (DFAs) to inaccurately describe ions. This method generally uses a Hartree-Fock electron density as input into the strongly constrained and appropriately normalised (SCAN) DFA. This method has shown promising results recently, including for aqueous ionic systems where it has been carefully validated in comparison to higher levels of quantum chemical theory.[52, 53, 54, 55, 56, 57] That said, it is likely that this method benefits from some degree of favourable cancellation of errors.[58]
@@
Here, we demonstrate that DC-DFT can be combined with NNPs to run accurate all-atom molecular dynamics of three electrolyte solutions, with explicit long-range electrostatics described by a continuum solvent model. These simulations require minimal training data (hundreds of frames) at a single concentration but can be used to simulate at both higher and lower concentration. We observe the formation of previously unknown Li cation dimers and almost identical anion pairing of chloride and bromide anions. We also demonstrate excellent agreement with experimental structural, kinetic, and thermodynamic properties. Secondly, we show that NNPs can be recursively trained on their own output to enable first principles coarse-grained continuum solvent molecular dynamics where only the ions are included. These simulations are dramatically faster and surprisingly can simulate complex, out-of-distribution behaviour such as crystal dissolution, and infinite dilution pairing free energies despite being trained only on moderate concentration solutions. This work therefore provides strong evidence against the notion that neural networks are only capable of interpolating on their training data. This workflow is outlined in Figure 1.
@@ 2 Results and Discussion
2.1 Low training data requirements
The training data for the LiCl, KCl and LiBr, all-atom NNP, trained with NequIP,[14] consists of a strikingly small training dataset of only 500-700 frames each extracted from MD simulation computed at the DC-r2SCAN level of theory with CP2K.[59, 60, 57] Each frame contain 4 cations, 4 anions and 80 water molecules corresponding to a 2.5M concentration. The faster r2SCAN level of theory was used to generate the initial trajectory, which was then resampled with DC-r2SCAN.
The total computational cost of generating this dataset was very reasonable, on the order of tens of thousands of CPU hours. It is therefore feasible to scale this approach to many different electrolyte solutions and conditions to build a large database of properties. Additional details on the training data generation process are outlined in the computational details section below.
@@ 2.1 Low training data requirements
Figure 2: (a) DFT forces are compared with the predictions of the NNP for LiCl. (b) Comparison of the ion-oxygen RDFs predictions of two NNPs trained with different seeds. Good agreement with neutron and X-ray diffraction (XRD) measurements of ion oxygen RDF is also observed. (c) Comparison of the cation-anion RDFs predictions of two NNPs trained with different initial seeds. (d) Comparison of the predicted O-O RDF from a pure water NNP-MD simulation with experimental XRD measurements[61] and the MB-pol water model.[62]
2.2 Long, large, and accurate simulations
@@
Two NNPs for each of LiCl, KCl, and LiBr were trained (NNP1 and NNP2) on the DC-r2SCAN dataset using different random seeds for the weight initialization and for splitting the data into training and validation sets. For LiCl, 200 frames were extracted from the NNP-MD simulation, and forces were recomputed with DFT for validation. Figure 2(a) demonstrates that the NNP very accurately estimates the forces during NNP-MD with an RMSE below 10 meVÅ-1.
@@
Parallel simulations were then run for over a nanosecond each with both NNP1 and NNP2 on a system six times larger than the original, containing 512 water molecules and 48 ions. Accessing this time and spatial scale is entirely infeasible with direct FPMD, yet it is critical for studying ion-ion interactions, which display long-range structure larger than the cell size used for the FPMD. The NNP-MD generates approximately 200 ps per day for the larger system on a single V100 GPU.
@@
Figure 2 compares the ion-solvent and ion-ion RDFs computed with the two separate NNPs, showing excellent agreement and demonstrating convergence and the reproducibility of the method. The ion-oxygen peak positions of 1.99, 2.78, 3.17, and 3.34 Å are in good agreement with neutron/X-ray diffraction measurements of 1.96, 2.79, 3.2, and 3.35 Å for lithium, potassium, chloride, and bromide, respectively.[63, 64] We use neutron diffraction measurements only for lithium due to the distorting effect of polarization on XRD measurements of its structure.[64] The ion-ion RDFs are also reproducible. Figure 2(c) demonstrates the strong specificity of cation-anion pairing.
@@
For systems of this size nonphysical artifacts can be observed every few nanoseconds, i.e., the formation of a very close-contact ion-ion pair. This indicates that we are operating at the minimum limit of training data. These errors can be corrected with various strategies such as additional resampling or active learning for generating new training data or with enhanced sampling or higher temperature sampling. These artifacts do not substantially alter the results of this study.
Water-water interactions help to determine the ion-ion interactions; it is therefore important to test the NNPs’ description of the pure water interactions. To do so we, can run a pure water simulation with the NNP trained on 2.5 M. Remarkably, Figure 2(d) shows essentially perfect agreement with experimental XRD measurements. The small difference in the first peak height is attributable to the neglect of quantum nuclear effects, which is verified by comparison them with MB-Pol, a state-of-the-art water model.[62]
The fact that, without any data on pure water, we arrive at a model that gives such accurate structural predictions is particularly promising and demonstrates the generalisability of this approach.
@@ 2.3 Lithium dimer formation
A surprising revelation from the simulations was the formation of lithium cation dimers, as shown in Figure 3, where one lithium penetrates into the first solvation layer around another lithium ion. The first solvation layer is defined by the first minimum in the Li-O radial distribution function (RDF), which is 2.8 Å. The smallest separation of the lithium ions observed in the simulations (2.68 Å) matches the separation of neutral covalent dilithium (2.67 Å) and is much smaller than the Li-Li distance in LiCl crystal (3.62 Å). The formation of this species is particularly surprising, given that this ion is an archetypal kosmotrope or water-structuring ion. Physically, this was believed to correspond to the formation of a tightly bonded first solvation layer of water molecules, which was thought to be impenetrable to other ions.[65]
@@
This counterintuitive finding may have important implications for many biological and chemical systems where lithium plays a critical role.[28] The transient nature of this pair means that this species could not feasibly be identified with direct FPMD simulation, whereas CMD simulations show no indication of it.[66] This is likely attributable to the neglect of charge transfer and polarization effects, which significantly mitigate the electrostatic repulsion. We have confirmed this effect is not an artifact of the NNP simulations by comparing the NNP predictions with the forces on the lithium dimer at the DFT level (Figure 12).
Figure 3: Lithium-oxygen and lithium-lithium RDFs demonstrate the formation of lithium cation dimers and an example snapshot from the MD simulation. .
2.4 Identical anion pairing
@@
A second surprising new finding is that the pairing of bromide ions in solution appears to be almost identical to that of chloride anions, as shown in Figure 4(a). This indicates that the pairing of anions in water is strongly determined by the surrounding water structure rather than by the inherent size of the anions. This is most likely related to the water-bridging structure formed by a water molecule forming a hydrogen bond with both anions simultaneously. This behavior is also totally different from the cation-cation pairing, where no bridging occurs and a significant difference between cations is observed, as shown in Figure 4(b). Figure 4(b) also demonstrates that Li pairing occurs similarly in LiBr electrolyte, indicating the anion is unlikely to play a critical role.
(a) Anion-Anion RDFs at 2.5 M
(b) Cation-cation RDFs at 2.5 M
Figure 4: Comparison of cation-cation and anion-anion RDFs for the electrolytes studied.
2.5 Thermodynamic and kinetic properties

### v4

@@ Abstract
Electrolyte solutions play critical role in a vast range of important applications, yet an accurate and scalable method of predicting their properties without fitting to experiment has remained out of reach, despite over a century of effort. Here, we combine state-of-the-art density functional theory and equivariant neural network potentials to demonstrate this capability, reproducing key structural, thermodynamic, and kinetic properties. We show that neural network potentials (NNPs) can be recursively trained on a subset of their own output to enable coarse-grained/continuum-solvent molecular simulations that can access much longer timescales than possible with all atom simulations. We observe the surprising formation of Li cation dimers along with identical anion-anion pairing of chloride and bromide anions. Finally, we reproduce simulate the crystal phase and infinite dilution pairing free energies despite being trained only on moderate concentration solutions. This approach should be scaled to build a greatly expanded database of electrolyte solution properties than currently exists.
@@ 1 Introduction
Molecular-scale processes, occurring at the level of thousands of atoms, are at the heart of chemistry and biology. These processes obey the laws of quantum mechanics and dictate the behavior of a vast range of crucial systems, yet they remain elusive, hidden from direct observation. We rely on indirect experiments and models to piece together their mysteries, but this limits our mastery over vital biological, chemical, and material systems. Unlocking these secrets could revolutionise our understanding and control of the world at its most fundamental level.
@@
Accurate and efficient first principles molecular dynamics simulations (FPMD) of these processes would be a transformatively useful tool for achieving this goal. This would enable the direct observation of key processes; the calculation of important properties such as free energies; and the generation of abundant training data for machine learning models. However, achieving this goal requires the solution to two inter-related problems: how to run calculations for sufficiently long times whilst employing sufficiently accurate models for the quantum mechanical interactions.
@@
A solution to the first problem is now possible thanks to neural network potentials (NNPs).[1, 2, 3, 4] This approach trains a neural network to reproduce quantum chemistry calculations, usually generated with density functional theory (DFT).[5, 6, 7, 8, 9, 10, 11, 12] NNPs are several orders of magnitude faster than direct calculations and are rapidly improving thanks to advances in the treatment of electrostatics.[13, 14, 7, 15, 16] and the use of equivariance,[17, 18, 19, 20, 21] which means that when the input to the NNP undergoes a symmetry transformation (like rotation or translation), the network’s output transforms in a predictable and consistent way under the same transformation.
Furthermore, coarse grained NNPs can also be trained and used to run coarse-grained molecular dynamics simulations that ignore irrelevant parts of a systems, enabling additional orders of magnitude acceleration.  [22, 23, 24, 25, 26] The simplest example of a coarse grained NNP is a continuum solvent NNP. This model takes as input the positions of all of the solutes in the system and outputs the average forces on them, where the averaging is over solvent configurations at equilibrium. The training data can therefore be obtained from equilibrium all atom MD simulation where the data on the solvent molecules is simply discarded. If the NNP correctly reproduces these average forces then it must be correctly modeling the free energy of the solutes as this corresponds to the ‘potential of mean force’.
So far these coarse grained NNPs have only been trained on classical all atom MD simulations, this is because DFT is too expensive to obtain the large datasets required. However, it should be possible to train them on all atom NNP-MD trained on DFT data.
Electrolyte solutions are a natural initial system to apply these tools to because determining the properties of electrolytes from first principles is a foundational problem of physical chemistry. Additionally, these solutions play a key role in a vast range of important processes and systems. Electrolytes are ubiquitous throughout chemical engineering and their equilibrium constants, solubilities, and reaction rates etc are critically important. The standard way to model these solutions today is with the century old Debye-Hückel theory augmented with empirically fitted parameters. Bypassing this reliance on empirical data [27] is of enormous practical value as obtaining high quality experimental data can be prohibitively challenging for many systems.
@@
However, studying electrolyte solutions exposes the second problem – how to achieve sufficient accuracy. Initial work on applying NNPs to this problem has shown significant promise. [28, 29, 13, 30, 2, 31, 32, 34, 15, 35, 36, 37, 38, 16, 39, 40, 41, 42, 43, 44, 45, 46] However, classical forcefields and low level DFT, such as generalised gradient approximation, are not sufficiently accurate [47, 48, 49, 50] to predict the complex dynamic interactions that govern electrolyte solution behaviour. Hence, quantitative prediction of important electrolyte solution properties such as activity coefficients, that depend sensitively on ion pairing has yet to be demonstrated. It has been particularly difficult to simulate ion pairing at low concentrations with NNP-MD because these pairs are much rarer than ion-solvent or solvent-solvent interactions and their energies are sensitive to the accuracy of the underlying quantum model, which means it is challenging to generate enough data with a high enough level of theory to train an accurate NNP.
@@
Figure 1: Workflow: Small, short first principles molecular dynamics (FPMD) simulations are run with CP2K. Energies/forces computed with DC-r2SCAN are used to train a neural network potential (NNP), which enables much faster simulations with a larger cell, while maintaining high accuracy. Forces and coordinates of the ions alone are output from the NNP-MD and used to train a coarse-grained/continuum-solvent NNP, which enables even faster coarse grained MD simulations to be performed and the simulation of crystals, despite being trained only on the solution-phase.
@@
Density-corrected DFT[51] (DC-DFT) offers a solution to the accuracy problem, as it has been shown to reduce errors, such as delocalization and self-interaction, which cause standard density functional approximations (DFAs) to inaccurately describe ions. DC-r2SCAN, which uses a Hartree-Fock electron density as input into the renormalized strongly constrained and appropriately normalised (SCAN) DFA, has demonstrated significant promise for aqueous ionic systems where it has been carefully validated in comparison to higher levels of quantum chemical theory; [52, 53, 54, 51, 55, 56] although favorable cancellation of errors may still be playing a role. [57] However, this accuracy comes at the expense of unfavourable computational cost, which hampers modelling of large systems or long time scales.
@@
Here, we demonstrate that DC-DFT can be used to train all atom NNPs to run accurate all-atom molecular dynamics of three electrolyte solutions, with explicit long-range electrostatics described by a continuum-solvent model. These all atom NNP-MD simulations require minimal training data (hundreds of frames) at a single concentration, allowing us to fit the NNP to accurate (but slow) DC-r2SCAN data, but can be used to simulate at both higher and lower concentration. We observe the formation of previously unknown Li cation dimers and almost identical anion pairing of chloride and bromide anions. We also demonstrate excellent agreement with experimental structural, kinetic, and thermodynamic properties.
We then demonstrate that the all atom NNP-MD simulations can be used to train coarse-grained/continuum-solvent NNPs where only the ions are included. These coarse grained NNPs can be used to run MD simulations that are dramatically faster, yet remarkably, they demonstrate complex, out-of-distribution behaviour such as crystal dissolution, and accurate infinite dilution pairing free energies; despite being trained only on moderate concentration solutions. This work thereby demonstrates the ability of NNPs to predict properties not directly seen in their training data, by exploiting fundamental physics like equivariance together with diverse and accurate quantum mechanical data from DC-r2SCAN. This workflow is outlined in Figure 1 and fully described in the Computational details section.
@@ 2 Results and Discussion
Figure 2: (a) DFT forces are compared with the predictions of the NNP for LiCl. (b) Comparison of the ion-oxygen RDFs predictions of two NNPs trained with different seeds. Good agreement with neutron and X-ray diffraction (XRD) measurements of ion oxygen RDF is also observed. (c) Comparison of the cation-anion RDFs predictions of two NNPs trained with different initial seeds. (d) Comparison of the predicted O-O RDF from a pure water all atom NNP-MD simulation with experimental XRD measurements[58] and the MB-pol water model.[59]
The NNP (built with NequIP[18]) exploits fundamental symmetries to build highly accurate potentials from strikingly small MD simulation data sets of 500-700 frames made up of 88 molecules each at only a single concentration (2.5 M). Its data efficiency enables the model to be trained on highly accurate yet costly DC-r2SCAN DFT calculations carried out with CP2K. [60, 61, 56] The resulting NNPs provide highly accurate and reproducible prediction of forces, ion-solvent, ion-ion and pure water structure as shown in Figure 2.
@@
Two NNPs were generated for each electrolyte using different random seeds, in order to validate them. For LiCl, 200 frames were extracted from the all atom NNP-MD simulation, and forces were recomputed with DFT for validation as shown in Figure 2(a) giving very accurate estimates the forces during NNP-MD with an RMSE below 10 meVÅ-1.
@@
Parallel all atom NNP-MD simulations were run for over a nanosecond each with both NNP1 and NNP2 on a system six times larger than the original, containing 512 water molecules and 48 ions. Accessing this time and spatial scale is entirely infeasible with direct FPMD, yet it is critical for studying ion-ion interactions, which display long-range structure larger than the cell size used for the FPMD. The all atom NNP-MD generates approximately 200 ps per day for the larger system on a single V100 GPU. In contrast the coarse grained simulations can produce 300 ps on a single Intel Xeon Scalable ‘Cascade Lake’ processors. Additionally the coarse graining accelerates the dynamics providing an additional speedup.
@@
Figure 2(b) and Figure 2(c) compares the ion-solvent and ion-ion RDFs computed with the two separate NNPs, showing excellent agreement and demonstrating convergence and the reproducibility of the method. The ion-oxygen peak positions of 1.99, 2.78, 3.17, and 3.34 Å are in good agreement with neutron/X-ray diffraction measurements of 1.96, 2.79, 3.2, and 3.35 Å for lithium, potassium, chloride, and bromide, respectively.[62, 63] The ion-oxygen peaks are also in good agreement with direct first principles simulation.[64] We use neutron diffraction measurements only for lithium due to the distorting effect of polarization on XRD measurements of its structure.[63] The ion-ion RDFs are also reproducible. Figure 2(c) demonstrates the strong specificity of cation-anion pairing. Experimental data on the ion-ion RDFs is extremely limited due to the challenges associated with extracting this signal from the background solvent-solvent and ion-solvent structure.
@@
For systems of this size nonphysical artifacts can be observed every few nanoseconds, i.e., the formation of a very close-contact ion-ion pair. This indicates that we are operating at the minimum limit of training data. These errors could be corrected with various strategies such as additional resampling or active learning for generating new training data or with enhanced sampling or higher temperature sampling. These artifacts do not substantially alter the results of this study.
Water-water interactions help to determine the ion-ion interactions; it is therefore important to test the NNPs’ description of the pure water interactions. To do so we, can run a pure water simulation with the all atom NNP trained on 2.5 M. Remarkably, Figure 2(d) shows essentially perfect agreement with experimental XRD measurements. “The small deviation from experiment in the first peak height can be attributed to the neglect of nuclear quantum effects (NQEs ), which modify the nuclear dynamics but not the electronic potential energy surface. We can verify this by comparing it to MB-Pol,[59] a water model which does reproduce the correct peak height when NQEs are included, but which gives very similar results to our model when they are neglected, suggesting that including NQEs in our simulation should also reduce the peak height. Excellent agreement between MB-pol and our model for the OH and HH RDFs is also shown in Figure S9. The fact that, without any data on pure water, we arrive at a model that gives such accurate structural predictions is particularly promising and demonstrates the generalisability of this approach.
(a) Lithium-oxygen and lithium-lithium RDFs showing the lithium dimer with an MD snapshot. The zoomed out full RDFs are shown in the inset.
(b) Cation-cation RDFs at 2.5 M
(c) Anion-Anion RDFs at 2.5 M
Figure 3: Cation-cation and anion-anions RDFs revealing the formation of lithium cation dimers and the almost identical pairing of bromide and chloride anions.
@@ 2.1 Lithium dimer formation
These all atom NNP-MD simulations immediately enable surprising new observations. Firstly, we observed the formation of lithium cation dimers, as shown in Figure 3(a), where one lithium penetrates into the first solvation layer around another lithium ion. The first solvation layer is defined by the first minimum in the Li-O radial distribution function (RDF), which is 2.8 Å. The smallest separation of the lithium ions observed in the all atom NNP-MD simulations (2.68 Å) is very similar to the separation of neutral covalent dilithium (2.67 Å) and is much smaller than the Li-Li distance in LiCl crystal (3.62 Å). The formation of this species is particularly surprising, given that lithium is an archetypal strongly hydrated ion. Physically, this was believed to correspond to the formation of a tightly bonded first solvation layer of water molecules, which was thought to be impenetrable to other ions.[65]
We can compute the coordination number of Li-Li coordination nubmer by integrating the RDF up to 4 Å  which is distance to the first minima of the Li-Li RDF. This gives a value of 0.05 at 2.4 M, indicating that there is a 5 % chance of finding a lithium ion dimer at 2.4 M.
@@
This counter intuitive finding may have important implications for many biological and chemical systems where lithium plays a critical role.[66] The transient nature of this pair means that this species could not feasibly be identified with direct FPMD simulation, whereas CMD simulations show no indication of it.[67] This is likely attributable to the neglect of charge transfer and polarization effects, which significantly mitigate the electrostatic repulsion, by transferring charge to the surrounding water molecules. Substantial charge transfer is a common observation of ions in water.[68] We have confirmed this effect is not an artifact of the all atom NNP-MD simulations by comparing the with DFT calculations on the lithium dimer structure (Figure S2). Figure 3(b) also demonstrates that Li pairing occurs similarly in LiBr electrolyte, indicating the anion is unlikely to play a critical role.
2.2 Water structure dominated anion pairing
@@
A second surprising new finding is that the pairing of bromide anions in solution appears to be almost identical to that of chloride anions, as shown in Figure 3(c). This indicates that the pairing of anions in water is strongly determined by the surrounding water structure rather than by the inherent size of the anions. This is likely related to the water-bridging structure formed by a water molecule that hydrogen bonds with both anions simultaneously. This behavior is also totally different from the cation-cation pairing, where no bridging occurs and a significant difference between cations is observed, as shown in Figure 3(b).
2.3 Ion pairing: prediction and validation
The reliability of the ion-ion RDFs can be validated by computing thermodynamic properties. Specifically, the activity coefficients as a function of concentration have been shown to be highly sensitive to the strength of ion-ion interactions. The accurate prediction of these quantities at very low concentration provided the original validation of Debye-Hückel theory.
Here, we use Kirkwood-Buff theory to compute a derivative of the log activities, achieving good experimental agreement as shown in Figure 4(a). The stronger pairing of KCl compared to LiCl is well reproduced. This is a classic example of a counterintuitive ‘specific ion effect’ induced by the solvent, which is reversed compared with the vacuum pairing strength. In contrast, the stronger pairing of LiCl compared with LiBr does follow the expected behaviour in vacuum, and is also well reproduced. Accurately reproducing these binding strengths is key to understanding many important phenomena such as the effect of ions on protein stability, i.e., the so-called ‘Hofmeister effect’.[69]